± Gravitational Force of Three Identical Masses Part A What is the magnitude of the net gravitational force Fgrav on the mass at the origin due to the other two Three identical very dense masses of 6500 kg each are placed on the x axis. One mass is at Zl = 120 cm, one is at the origin, and one is at r2- 330 cm masses? Take the gravitational constant to be G 6.67x10-11 N m2/kg Express your answer in newtons to three significant figures Hints grav Submit My Answers Give Up Incorrect; Try Again; 5 attempts remaining
Answer #1

m1 = 6500 kg

m2 = 6500 kg


m3 = 6500 kg


x1 = 0

x2 = -120 cm = -1.2 m

x3 = 330 cm = 3.3 m


force exerted by m2 on m1 F21x = -G*m1*m2/x1^2

F21x = -6.67*10^-11*6500*6500/1.2^2

F21x = -0.00195 N

force exerted by m3 on m1 F31x = -G*m1*m3/x1^2

F31x = 6.67*10^-11*6500*6500/3.3^2

F31x = 0.00026 N

Fgrav = F21x + F31x

Fgrav = -0.00195 + 0.00026


Fgrav = -0.00169 N